A helium-filled weather balloon has a volume of 595 L at 14.9°C and 751 mmHg. It is released and rises to an altitude of 5.55 km, where the pressure is 451 mmHg and the temperature is –17.1°C. The volume of the balloon at this altitude is ___ L.
Answer #1

T °C +213.15 Initial conditims. v=595L T = (14.9+273.15 JK T = 288.05k P-7 Si mmHg altitude of 5.55km Final conditions V = ?